37000 people at ball game. Receipts were $248,000. How many people paid $12 for reserved seats, and how many paid $5 for general admission?
Answer by checkley77(12569)

12x+5(37,000-x)=248,000
12x+185,000-5x=248,000
7x=248,000-185,000
7x=63,000
x=63,000/7
x=9,000 is the number of $12 tickets sold.
37,000-9,000=28,000 number of $5 tickets sold.
Proof:
12*9,000+5*28,000=248,000
108,000+140,000=248,000
248,000=248,000